1. A 60 kg woman walking through an alley accidentally broke the heel of her left high-heeled



shoes. Now the woman tries to balance on one heel of a pair of high-heeled shoes. If the heel is circular and has a radius of 0.500 cm, what pressure does she exert on the floor?

1
Expert's answer
2022-06-10T18:32:50-0400

"p=\frac FS=\frac{mg}{\pi r^2}=7 .5~MPa."